# Improved Search-to-Decision Reduction for Random Local Functions

### v1

@@ Abstract
We present a new search-to-decision reduction for random local functions defined by any predicate of constant arity. Given any efficient algorithm that can distinguish, with advantage $\varepsilon$, the output of a random local function with $m$ outputs and $n$ inputs from random, our reduction produces an efficient algorithm that can invert such functions with $\tilde{O}(m(n/\varepsilon)^{2})$ outputs, succeeding with probability $\Omega(\varepsilon)$. This implies that if a family of local functions is one-way, then a related family with shorter output length is family of pseudo-random generators.
@@ 1 Introduction
To be more precise, this is a function $f_{G,P}:\mathbb{F}_{2}^{n}\rightarrow\mathbb{F}_{2}^{m}$ such that on an input $s\in\mathbb{F}_{2}^{n}$, the $i$-th output bit of the function $f_{G,P}(s)_{i}$ is $P(s_{j_{1}},\dots,s_{j_{d}})$, where each $j_{k}\in[n]$ is a randomly selected index. The collection of indices for each output bit is represented as a tuple (or a hyperedge) $S_{i}=(j_{1},\dots,j_{d})\in[n]^{d}$. The collection of tuples is described as a hypergraph $G=(S_{i})_{i\in[m]}$. Goldreich conjectured that if the predicate $P$ is appropriately chosen and $m$ is not too large or small, then such a function is one-way with high probability. [^1] Goldreich originally studied a deterministic version of these functions where the hypergraph is fixed to be one with certain expansion properties. Nevertheless, as random graphs are expanding with high probability, his conjectures also imply the one-wayness of the functions we describe here.
Crucial to the hardness of inverting these functions is the choice of the predicate $P$ and the relationship between the output and input lengths $m$ and $n$. For instance, it is known that, regardless of the choice of predicate $P$, inversion can be performed efficiently for some $m=O(n^{\frac{1}{2}\left\lfloor 2d/3\right\rfloor}\log n)$ [App16, Corollary 3.6]. Over the past few decades, there has been an active line of work studying the complexity of inversion for various predicates and parameter settings [Gol11, AHI05, BQ09, CEMT09, Its10, OW14, AL16, App16, CDM+18, COST19, YGJL22, DMR23]. We discuss this work in more detail in Section 1.3.
@@
While conceptually simple, the problem of inverting a random local function is related to several important problems from various areas of Computer Science. Viewing the outputs of the function as constraints that the input needs to satisfy, this is an instance of a random constraint satisfaction problem (CSP) [AIK08, CEMT09, RRS17, GHKM23, DV21].
Furthermore, suppose the predicate $P$ is a noisy XOR function, i.e. $P$ computes the parity of its inputs and flips it with a small probability. This corresponds to the sparse learning parity with noise problem (LPN) [Fei02, Ale03, FGKP06, CSZ24, BBTV24], which has many applications in public key cryptography [ABW10], machine learning [BEG+22], indistinguishability obfuscation [RVV24] and more [CGHKV24, DJ24, ABCM25, BCM+25]. If we then interpret the hypergraph as describing the generator matrix of a linear code, inverting the local function corresponds to decoding its noisy codewords from random noise. Similar to how Linear Regression is usually interpreted, this may also be seen as a learning problem where the task is to learn the hidden linear function $\langle x,\cdot\rangle$ given many noisy evaluations.
@@
Interpreting the problem of inversion as a search problem, one can naturally define the corresponding decision problem as distinguishing the output of a random local function $(G,f_{G,P}(s))$, on a random input $s$, from $(G,b)$, where $b\in\mathbb{F}_{2}^{m}$ is a uniformly random and independent of $G$. Viewing the input $s$ as the random seed, the hardness of the decision problem would imply that the family of local functions is a family of Pseudo-Random Generators (PRG).
Given their locality, Such PRG’s could be very efficient to compute, but have a couple of limitations. For example, with probability $1/\mathrm{poly}(n)$, the graph $G$ has a pair of repeated hyperedges, in which case the corresponding outputs bits of $f_{G,P}$ can be used to distinguish it from random. The best bound on an adversary’s distinguishing advantage we can hope for against a random local function is thus some $1/\mathrm{poly}(n)$; we refer to such PRG’s as “weak” PRG’s. Second, known algorithms place a limit of $\tilde{O}(n^{\left\lfloor 2d/3\right\rfloor/2})$ on the stretch of this PRG if a $d$-ary predicate is used [App16], and even stronger limits are known for certain classes of predicates [BQ09, AL16].
Regardless, it is still possible to construct a local PRG with negligible distinguishing advantage (a “strong” PRG) from a random local function by processing the output further. Applebaum [App12] showed how to get a strong local PRG with linear stretch by applying a local extractor to the output of an unpredictable random local function. Later, Applebaum and Kachlon [AK19] effectively resolved these issues by showing that any weak local PRG that has non-trivial polynomial stretch can be transformed into a strong local PRG with any desired polynomial stretch, at the cost of some degradation of the locality.
@@
Following this discussion, if we can reduce the above search problem to the decision problem, we can construct local PRG’s from the one-wayness of random local functions. Further, given the fundamental nature of the family of local functions and the aforementioned connections to CSPs, decoding, etc., such reductions would be of broad interest even outside the context of local cryptography.
@@
They extend the above result to accommodate noisy predicates $P$, incurring an additional $\log n$ factor in the output length required by the search algorithm [App12, Theorem 4.5]. They also show that if some output bit of a random local function is predictable with constant advantage given all previous output bits, then the function can be inverted with constant probability as well; this holds without any blowup in output length and without requiring the predicate to be sensitive, but falls short of leading to pseudo-randomness or a reduction to the decision problem.
@@
  • [BSV19] showed that there exists an algorithm for the search problem with $O(m(m/\varepsilon)^{2/d})$ outputs, maintaining the same locality constant $d$, but succeeding only with small probability $\exp(-\tilde{O}(m/\varepsilon)^{6/d})$.
  • [BRT25] showed that there exists an algorithm for the search problem with $O(nm+n/\eta^{2}\varepsilon^{2(d-2)})$ outputs and $n$ inputs for the noisy $(d+1)$-ary XOR predicate. They further generalise their result to any sensitive predicate, though again the predicates in the decision and search problems are slightly different.
@@ 1.1 Our Contribution
The main improvement of our work over all existing search-to-decision reductions is the removal of the requirement of the predicate being sensitive. This generalisation opens up the possibility of constructing local PRG’s (including strong ones, following [AK19]) using the hardness of inverting a wider range of local functions. Lack of sensitivity is potentially beneficial from a cryptographic perspective, as it means one less form of structure that might enable algorithms. This aligns with the results of [DMR23], who proposed alternatives to the XOR-MAJ predicate for use in random local functions. For instance, they identified a predicate on 9 variables that has maximal algebraic immunity and resilience among all such predicates (both being properties needed for security), and this predicate is non-sensitive.
@@
Our analysis depends only minimally on the predicate itself, relying instead on the mixing properties of certain transformations we perform on the hypergraph. We hope that these techniques will be useful in similar reductions beyond the context of local functions as well.
@@ 1.2 Overview of Techniques
We will first provide more details on the construction of the predictors. Suppose we have a decision algorithm $\mathsf{D}$ that has advantage $\varepsilon$. Our goal will be to construct a set of algorithms $\mathsf{S}_{i}$, such that for a large enough fraction of secrets $s\in\mathbb{F}_{2}^{n}$, there are a pair of numbers $eq_{s},neq_{s}\in[0,1]$ where $eq_{s}>neq_{s}$, such that the following holds for all $i\in[2,n]$:
@@
On the other hand, when $s_{a}\neq s_{b}$, then in general $f_{G,P}(s)$ is not the same as $f_{T_{a,b}(G),P}(s)$, and the relationship between $T(G)$ and $f_{G,P}(S)$ becomes less coupled that that between $G$ and $f_{G,P}(s)$. In fact, we can show that after we apply about $t=O(n\log(nm/\varepsilon))$ such transformations, the transformed hypergraph will become independent of $f_{G,P}(s)$.
Precisely, we show that for any hypergraph $G$, the distribution of $T_{a_{t},b_{t}}\circ\ldots\circ T_{a_{1},b_{1}}(G)$ is close to uniformly sampling from $G_{n,m,d}$. One can think of the transformation as a Markov process, and $O(n\log(nm/\varepsilon))$ is the mixing time of this process. [^2] The proof of this claim is inspired by the convergence of randomised gossip algorithms [BGPS06]. This implies that with an independently sampled $G^{\prime}\leftarrow G_{n,m,d}$,
@@
To see why this might true, note that after about $t=\Omega(n\log(nm/\varepsilon))$ such randomly chosen transformations, every vertex in every hyperedge of $G$ would have been touched by at least a few of the transformations. As the random assignment in $T_{a,b}$ happens independently for each $a$ or $b$ that appears as a vertex, this process quickly randomises the whole hypergraph.
@@ 1.3 Related Work
Goldreich [Gol11] originally proposed using a randomly chosen predicate, while cautioning against linear, degenerate, or otherwise structured predicates that lead to easily solvable equation systems. It is known that myopic and drunken backtracking algorithms (algorithms that only look at the instance locally each time they make a decision; this includes many powerful SAT solvers) do not perform well on predicates with some linear component $P(x_{1},\dots,x_{d})=x_{1}+\dots+x_{k}+Q(x_{k+1},\dots,x_{d})$ for an arbitrary $Q$ with $k>3d/4$ for some large enough $d$ on $n$ variables and $n$ outputs [AHI05, CEMT09, Its10]. For the XOR-AND3,2 predicate $P(x_{1},\dots x_{5})=x_{1}+x_{2}+x_{3}+x_{4}x_{5}$, [OW14] show pseudorandomness up to output length $n^{1.499}$ against $\mathbb{F}_{2}$-linear tests and a wide class of semi-definite programming algorithms. The work of [ABR12] showed a dichotomy – for output length $n^{1+\delta}$, every choice of the predicate results in the random local function that either is secure against linear tests with high probability, or is insecure with high probability.
@@
[CDM+18] looked at the concrete security of Goldreich’s function and developed a sub-exponential time attack and analysed the efficiency of algebraic attacks, such as Gröbner basis methods. [YGJL22] built on their work, improving the time complexity. [COST19] showed that the assumption of the hypergraph being a good expander is not sufficient; the neighbour function of the graph must also have high circuit complexity.
@@
On attacks on local PRG’s, it is known that there exists a subexponential time algorithm for distinguishing superlinear stretch local PRG with noticeable probability [AIK06, Zic17, Üna23]. A related variant of local PRG is also studied by [ABCM25], where the input string is not uniformly sampled. Assuming the hardness of sparse LPN, they showed that superlinear stretch with negligible distinguishing advantage is possible.
@@ 2 Preliminaries
Notations.
Given a binary string $x$, we use $x_{i}$ to denote the $i$-th bit in the binary string. Given two distributions $D_{1},D_{2}$. If the two distributions are equivalent, we write $D_{1}\approx D_{2}$. If $x\leftarrow D_{1}$, this means $x$ is sampled from $D_{1}$. Here, $D_{1}$ can also be a set, in which case this denotes sampling uniformly from the set. We denote a finite field of size $p$ as $\mathbb{F}_{p}$ where $p$ is a prime. For any $\eta\in[0,1]$, we denote by $Bern(\eta)$ the Bernoulli distribution with parameter $\eta$.
@@
where $e$ is a Bernoulli random variable. A constant correlation also implies that the Bernoulli parameter is constant. An algorithm to solve the search problem is constructed by reducing the problem to a $c$-sparse noisy linear system. Having $m=\Omega(n^{c/2})$ gives roughly $\Omega(n)$ pairs of equations where c1c-1 variables are the same, while the rest of the variables are disjoint. Each pair then have its equations combined to obtain 2-sparsity noisy linear equations with constant error rate. Since there exists an efficient algorithm to solve a 2-sparsity noisy linear system with a constant error rate, we are done. ∎

### v2

@@ Abstract
We present a new search-to-decision reduction for random local functions defined by any predicate of constant arity. Given any efficient algorithm that can distinguish, with advantage $\varepsilon$, the output of a random local function with $m$ outputs and $n$ inputs from random, our reduction produces an efficient algorithm that can invert such functions with $\tilde{O}(m(n/\varepsilon)^{2})$ outputs, succeeding with probability $\Omega(\varepsilon)$. This implies that if a family of local functions is one-way, then a related family with shorter output length is a family of pseudo-random generators.
@@ 1 Introduction
To be more precise, this is a function $f_{G,P}:\mathbb{F}_{2}^{n}\rightarrow\mathbb{F}_{2}^{m}$ such that on an input $s\in\mathbb{F}_{2}^{n}$, the $i$-th output bit of the function $f_{G,P}(s)_{i}$ is $P(s_{j_{1}},\dots,s_{j_{d}})$, where each $j_{k}\in[n]$ is a randomly selected index. The collection of indices for each output bit is represented as a tuple (or an ordered hyperedge) $S_{i}=(j_{1},\dots,j_{d})\in[n]^{d}$. The collection of tuples is described as a hypergraph $G=(S_{i})_{i\in[m]}$. Goldreich conjectured that if the predicate $P$ is appropriately chosen and $m$ is not too large or small, then such a function is one-way with high probability. [^1] Goldreich originally studied a deterministic version of these functions where the hypergraph is fixed to be one with certain expansion properties. Nevertheless, as random graphs are expanding with high probability, his conjectures also imply the one-wayness of the functions we describe here.
Crucial to the hardness of inverting these functions is the choice of the predicate $P$ and the relationship between the output and input lengths $m$ and $n$. For instance, it is known that, regardless of the choice of predicate $P$, inversion can be performed efficiently for some $m=O(n^{\frac{1}{2}\left\lfloor 2d/3\right\rfloor}\log n)$ [APP16, Corollary 3.6]. Over the past few decades, there has been an active line of work studying the complexity of inversion for various predicates and parameter settings [GOL11, AHI05, BQ09, CEM+09, ITS10, OW14, AL16, APP16, CDM+18, CST19, YGJ+22, DMR23]. We discuss this work in more detail in Section 1.3.
@@
While conceptually simple, the problem of inverting a random local function is related to several important problems from various areas of Computer Science. Viewing the outputs of the function as constraints that the input needs to satisfy, this is an instance of a random constraint satisfaction problem (CSP) [AIK08, CEM+09, RRS17, GHK+23, DV21].
Furthermore, suppose the predicate $P$ is a noisy XOR function, i.e. $P$ computes the parity of its inputs and flips it with a small probability. This corresponds to the sparse learning parity with noise problem (LPN) [FEI02, ALE03, FGK+06, CSZ24, BBT+24], which has many applications in public key cryptography [ABW10], machine learning [BEG+22], indistinguishability obfuscation [RVV24] and more [CHK+24, DJ24, ABC+25, BCM+25]. If we then interpret the hypergraph as describing the generator matrix of a linear code, inverting the local function corresponds to decoding its noisy codewords from random noise. Similar to how Linear Regression is usually interpreted, this may also be seen as a learning problem where the task is to learn the hidden linear function $\langle x,\cdot\rangle$ given many noisy evaluations.
@@
Interpreting the problem of inversion as a search problem, one can naturally define the corresponding decision problem as distinguishing the output of a random local function $(G,f_{G,P}(s))$, on a random input $s$, from $(G,b)$, where $b\in\mathbb{F}_{2}^{m}$ is uniformly random and independent of $G$. Viewing the input $s$ as the random seed, the hardness of the decision problem would imply that the family of local functions is a family of Pseudo-Random Generators (PRG).
Given their locality, such PRGs could be very efficient to compute, but have a couple of limitations. For example, with probability $1/\mathrm{poly}(n)$, the graph $G$ has a pair of repeated hyperedges, in which case the corresponding output bits of $f_{G,P}$ can be used to distinguish it from random. The best bound on an adversary’s distinguishing advantage we can hope for against a random local function is thus some $1/\mathrm{poly}(n)$; we refer to such PRGs as “weak” PRGs. Second, known algorithms place a limit of $\tilde{O}(n^{\left\lfloor 2d/3\right\rfloor/2})$ on the stretch of this PRG if a $d$-ary predicate is used [APP16], and even stronger limits are known for certain classes of predicates [BQ09, AL16].
Regardless, it is still possible to construct a local PRG with negligible distinguishing advantage (a “strong” PRG) from a random local function by processing the output further. Applebaum [APP12] showed how to get a strong local PRG with linear stretch by applying a local extractor to the output of an unpredictable random local function. Later, Applebaum and Kachlon [AK19] effectively resolved these issues by showing that any weak local PRG that has non-trivial polynomial stretch can be transformed into a strong local PRG with any desired polynomial stretch, at the cost of some degradation of the locality. The key idea is to apply an additional layer of a local randomness extractor to the output of the weak PRG.
Theorem 1.1 ([AK19, Theorem 2.12] weak-to-strong compiler).
For every constant $d\in N$, $a>0$ and $c,c^{\prime}>1$, there exists a constant $d^{\prime}$ for which the following holds. Any $d$-local PRGs that stretch $n$ bits into $n^{c}$ bits with at most $\varepsilon=1/n^{a}$ distinguishing advantage can be converted into $d^{\prime}$-local PRGs that stretch $n$ bits to $n^{c^{\prime}}$ bits with at most a negligible distinguishing advantage.
@@
Following this discussion, if we can reduce the above search problem to the decision problem, we can construct local PRGs from the one-wayness of random local functions. Further, given the fundamental nature of the family of local functions and the aforementioned connections to CSPs, decoding, etc., such reductions would be of broad interest even outside the context of local cryptography.
@@
They extend the above result to accommodate noisy predicates $P$, incurring an additional $\log n$ factor in the output length required by the search algorithm [APP12, Theorem 4.5]. The proof uses a distinguisher to derive a “next-bit predictor” that can predict an output bit given all previous output bits. It is then subsequently used to solve the search problem.
They also show that if the next-bit predictor gives a constant advantage, then the function can be inverted with constant probability; this holds without any blowup in output length and without requiring the predicate to be sensitive, but falls short of leading to pseudo-randomness or a reduction to the decision problem.
@@
  • [BSV19] showed that there exists an algorithm for the search problem with $O(m(m/\varepsilon)^{2/d})$ outputs, maintaining the same locality constant $d$, but succeeding only with small probability $\exp(-\tilde{O}(m/\varepsilon)^{6/d})$. Their approach relies on a new approximate local list-decoding algorithm for the $d$-XOR code at large distances.
  • [BRT25] showed that there exists an algorithm for the search problem with $O(nm+n/\eta^{2}\varepsilon^{2(d-2)})$ outputs and $n$ inputs for the noisy $(d+1)$-ary XOR predicate. They further generalise their result to any sensitive predicate, though again the predicates in the decision and search problems are slightly different. The main approach is a transformation that removes a sensitive variable from the predicate, creating a distributional shift. This shift then allows the construction of a predictor that solves the search problem.
@@ 1.1 Our Contribution
The main improvement of our work over all existing search-to-decision reductions is the removal of the requirement of the predicate being sensitive. This generalisation opens up the possibility of constructing local PRGs (including strong ones, following Theorem 1.1) using the hardness of inverting a wider range of local functions. Lack of sensitivity is potentially beneficial from a cryptographic perspective, as it means one less form of structure, which may make certain classes of attacks harder to mount.
@@
Our analysis depends only minimally on the predicate itself, relying instead on the mixing properties of certain transformations we perform on the hypergraph. We hope that these techniques will be useful in similar reductions beyond the context of local functions as well. We note that our reduction does not achieve the same level of sample complexity as those for sensitive predicates (approximately losing a factor of $n$), but we introduce techniques that we hope will allow future work to close the remaining gap between our loss factor and the tighter bounds known for sensitive predicates.
Optimal Reduction
For sensitive predicates, there is currently no proven generic sample complexity gap between the search problem and the decision problem. For a general predicate, however, there is a trivial gap of $\sqrt{n}$. Consider a random local function with $m$ outputs and $n$ inputs, with the predicate $P$ being the majority of 3 variables. One can solve the distinguishing task by finding a pair of outputs that depend on the same input. This is because for a majority predicate, sharing an input implies correlated output. With around $m=\sqrt{n}$ samples, such a pair appears with constant probability. However, it is information theoretically impossible to solve the search with just $\sqrt{n}$ outputs. Therefore, the best possible search to decision reduction reduces search with $\tilde{O}_{\varepsilon}(m\sqrt{n})$ samples to decision with $m$ samples. This shows that our reduction is $\tilde{O}_{\varepsilon}(n^{1.5})$ factor far from optimal.
@@ 1.2 Overview of Techniques
We will first provide more details on the construction of the predictors. Suppose we have a decision algorithm $\mathsf{D}$ that has advantage $\varepsilon$. Our goal will be to construct a set of algorithms $\mathsf{S}_{i}$, such that for a large enough fraction of secrets $s\in\mathbb{F}_{2}^{n}$, there is a pair of numbers $eq_{s},neq_{s}\in[0,1]$ where $eq_{s}>neq_{s}$, such that the following holds for all $i\in[2,n]$:
@@
On the other hand, when $s_{a}\neq s_{b}$, then in general $f_{G,P}(s)$ is not the same as $f_{T_{a,b}(G),P}(s)$, and the relationship between $T(G)$ and $f_{G,P}(S)$ becomes less coupled than between $G$ and $f_{G,P}(s)$. In fact, we can show that after we apply about $t=O(n\log(nm/\varepsilon))$ such transformations, the transformed hypergraph will become independent of $f_{G,P}(s)$.
Precisely, we show that for any hypergraph $G$, the distribution of $T_{a_{t},b_{t}}\circ\ldots\circ T_{a_{1},b_{1}}(G)$ is close to uniformly sampling from $G_{n,m,d}$. One can think of the transformation as a Markov process, and $O(n\log(nm/\varepsilon))$ is the mixing time of this process. [^2] The proof of this claim is inspired by the convergence of randomised gossip algorithms [BGP+06]. This implies that with an independently sampled $G^{\prime}\leftarrow G_{n,m,d}$,
@@
To see why this might be true, note that after about $t=\Omega(n\log(nm/\varepsilon))$ such randomly chosen transformations, every vertex in every hyperedge of $G$ would have been touched by at least a few of the transformations. As the random assignment in $T_{a,b}$ happens independently for each $a$ or $b$ that appears as a vertex, this process quickly randomises the whole hypergraph.
@@ 1.3 Related Work
Goldreich [GOL11] originally proposed using a randomly chosen predicate, while cautioning against linear, degenerate, or otherwise structured predicates that lead to easily solvable equation systems. It is known that myopic and drunken backtracking algorithms (algorithms that only look at the instance locally each time they make a decision; this includes many powerful SAT solvers) do not perform well on predicates with some linear component $P(x_{1},\dots,x_{d})=x_{1}+\dots+x_{k}+Q(x_{k+1},\dots,x_{d})$ for an arbitrary $Q$ with $k>3d/4$ for some large enough $d$ on $n$ variables and $n$ outputs [AHI05, CEM+09, ITS10]. For the XOR-AND3,2 predicate $P(x_{1},\dots x_{5})=x_{1}+x_{2}+x_{3}+x_{4}x_{5}$, [OW14] show pseudorandomness up to output length $n^{1.499}$ against $\mathbb{F}_{2}$-linear tests and a wide class of semi-definite programming algorithms. The work of [ABR12] showed a dichotomy – for output length $n^{1+\delta}$, every choice of the predicate results in the random local function that either is secure against linear tests with high probability, or is insecure with high probability.
@@
[CDM+18] looked at the concrete security of Goldreich’s function and developed a sub-exponential time attack and analysed the efficiency of algebraic attacks, such as Gröbner basis methods. [YGJ+22] built on their work, improving the time complexity. [CST19] showed that the assumption of the hypergraph being a good expander is not sufficient; the neighbour function of the graph must also have high circuit complexity.
@@
On attacking local PRGs, it is known that there exists a subexponential time algorithm for distinguishing superlinear stretch local PRGs with noticeable probability [AIK06, ZIC17, ÜNA23]. A related variant of local PRG is also studied by [ABC+25], where the input string is not uniformly sampled. Assuming the hardness of sparse LPN, they showed that superlinear stretch with negligible distinguishing advantage is possible.
@@ 2 Preliminaries
Notation.
Given a binary string $x$, we use $x_{i}$ to denote the $i$-th bit in the binary string. Given two distributions $D_{1},D_{2}$, If the two distributions are equivalent, we write $D_{1}\approx D_{2}$. If $x\leftarrow D_{1}$, this means $x$ is sampled from $D_{1}$. Here, $D_{1}$ can also be a set, in which case this denotes sampling uniformly from the set. We denote a finite field of size $p$ as $\mathbb{F}_{p}$ where $p$ is a prime. For any $\eta\in[0,1]$, we denote by $Bern(\eta)$ the Bernoulli distribution with parameter $\eta$.
@@
where $e$ is a Bernoulli random variable. A constant correlation also implies that the Bernoulli parameter is constant. An algorithm to solve the search problem is constructed by reducing the problem to a $c$-sparse noisy linear system. Having $m=\Omega(n^{c/2})$ gives roughly $\Omega(n)$ pairs of equations where c1c-1 variables are the same, while the rest of the variables are disjoint. Each pair then has its equations combined to obtain 2-sparsity noisy linear equations with constant error rate. Since there exists an efficient algorithm to solve a 2-sparsity noisy linear system with a constant error rate, we are done. ∎